OpenStudy (anonymous):

Solve this system of equations using substitution. a+b+c=180 b=3a c=5a b=?

OpenStudy (anonymous):

Have you tried it yet?

OpenStudy (anonymous):

need help please!!!

OpenStudy (anonymous):

How far did you get?

OpenStudy (anonymous):

thats eassy just substitute for a & c then get b

OpenStudy (anonymous):

i dont know where to start

OpenStudy (anonymous):

Plug in b=3a and c=5a into a+b+c=180. Solve this for a. When you know a, you can easily solve b=3a for b.

OpenStudy (anonymous):

can you show me the step for this question?

OpenStudy (anonymous):

Sure. Since b=3a and c=5a, when we substitute these values into a+b+c=180, we get a+3a+5a=180. So 9a+180. Therefore, a=20. Does that make sense? Now you plug the value for a into the equation b=3a. What do you get for b then?

OpenStudy (anonymous):

simply replace b = 3a and c = 5a to get b then you can continue from there easy enough

OpenStudy (anonymous):

is it b=60

OpenStudy (anonymous):

Yep. Good job.

OpenStudy (anonymous):

plug it in and see

OpenStudy (anonymous):

You can check your work by plugging in a=20, b=3a=3(20)=60, and c=5a=5(20)=100 into a+b+c=180 and making sure that it works (which it does).

OpenStudy (anonymous):

I need help on C=5a

OpenStudy (anonymous):

you have what a equals and you have what b equals put them together in your intial equation and check to see if it works.